Chip’s Home Brew Whiskey management forecasts that if the firm sells each bottle of Snake-Bite for $20, then the demand for the product will be 15,000 bottles per year. Whereas Sales will equal only 89 percent of this amount if the price is raised 10 percent. Chip’s variable cost per bottle is $10, and the total fixed cash cost for the year is $100,000. Depreciation and amortization charges are $20,000, and the firm has a 30 percent marginal tax rate. Management anticipates an increased working capital need of $3,000 for the year. What will be the effect of a 10 percent price increase on the firm’s FCF for the year?
Answer #1

Let's call the forecasts corresponding to the two situations as Forecast 1 and 2.

Under forecast 1:

  • Sale Price, P = $ 20 per bottle
  • Sale volume, Q = 15,000 bottles

Under forecast 2

  • Sale price, P = 10% more than forecast 1 sale price = $ 20 x (1 + 10%) = $ 22 per bottle
  • Sale Volume, Q = 89% of forecast 1 sale volume = 89% x 15,000 = 13,350 bottles

Now, please see the table below. The second column titled "Linkage" explains how each row has been calculated.

Parameter

Linkage

Forecast 1

Forecast 2

Sale Price

P

20

22

Sale Volume

Q

15,000

13,350

Revenue

R = P x Q

300,000

293,700

Variable cost per bottle

V

10

10

Total Variable Cost

VC = V x Q

150,000

133,500

Fixed Cost

FC

100,000

100,000

Depreciation & amortization expenses

D

20,000

20,000

Operating income

EBIT = R - VC - FC -D

30,000

40,200

Marginal tax rate

t

30%

30%

Taxes

T = t x EBIT

9,000

12,060

NOPAT

EBIT - T

21,000

28,140

Free Cash flow calculation

NOPAT = EBIT x (1 - t)

NOPAT

21,000

28,140

[+] Depreciation & amortization

D

20,000

20,000

[-] Increase in working capital

Increase in WC

3,000

3,000

FCFF

NOPAT + D - Increase in WC

38,000

45,140

Hence,  the effect of a 10 percent price increase on the firm’s FCF for the year = FCFF under Forecast 2 - FCFF under forecast 1 = 45,140 - 38,000 = $ 7,140 increase.